# Structure-Fair Quantum Circuit Complexity: An Auditable Information-Theoretic Lower Bound

### v2

@@ Appendix F Generalized Settings and Framework Robustness
This appendix aims to explore the applicability of the Reference-Contingent Complexity (RCC) framework in more generalized settings and to address profound questions regarding the objectivity of its core constructs. Section F.1 generalizes the theory to reference sets defined by non-Abelian symmetries. Section F.2 provides a rigorous scheme for a conservative lower bound for handling systems that require an energy or mode cutoff (such as quantum field theories). Finally, Section F.3 provides physical, objectivization criteria for the selection of the reference set and proves the robustness of the framework with respect to this choice[62, 63].
@@ F.1 Non-Abelian Reference Sets
$\Pi_{R}=\bigoplus_{\lambda\in\Lambda}(I_{V_{\lambda}}\otimes Q_{\lambda})\\ d_{R}=\mathrm{Tr}(\Pi_{R})=\sum_{\lambda\in\Lambda}d_{\lambda}\cdot r_{\lambda}\\ \sigma_{R}=\Pi_{R}/d_{R}$ (F.1)
@@ F.2 Conservative Lower Bound under Energy or Mode Cutoff
$D(\rho_{\mathrm{phys}}||\tilde{\sigma}_{R}^{(\Lambda)})\geq D_{\mathrm{Bern}}(q||1-\delta)+q\cdot D(\rho^{(\Lambda)}||\sigma_{R}^{(\Lambda)})$ (F.4)
@@
$D(\rho_{\mathrm{phys}}||\tilde{\sigma}_{R}^{(\Lambda)})\geq q\cdot D(\rho^{(\Lambda)}||\sigma_{R}^{(\Lambda)})$ (F.5)
@@ F.3 Acceptability and Objectivity of the Reference Set
$|C_{R}(\rho)-C_{R^{\prime}}(\rho)|\leq\left|\frac{\log(d_{R})}{\log(\Gamma_{R})}-\frac{\log(d_{R^{\prime}})}{\log(\Gamma_{R^{\prime}})}\right|+\left|\frac{1}{\log(\Gamma_{R})}-\frac{1}{\log(\Gamma_{R^{\prime}})}\right|\cdot S(\rho)$ (F.6)
@@
$0\leq C_{R}(\rho)-C_{R^{\prime}}(\rho)=\frac{\log(d_{R})-\log(d_{R^{\prime}})}{\log(\Gamma_{R})}$ (F.7)
@@ Appendix G Near-Tightness of Reference-Contingent Complexity
This appendix aims to provide constructive, non-trivial, positive evidence for Conjecture 6.1 (the near-tightness of $RCC$), which was proposed in Section 6.2 of the main text. The conjecture posits that Reference-Contingent Complexity $C_{R}(\rho)$ is not only a lower bound on the universal circuit complexity $C_{\mathrm{opt}}(\rho)$, but that for a class of well-structured families of states, it can also be bounded from above by $C_{R}(\rho)$. We will systematically corroborate this conjecture by constructing complexity upper bounds for three core families of states that are widely present in quantum information and physics: efficiently samplable classical diagonal states, low-rank states constructible via hierarchical projections, and efficiently encodable logical pure states with sparse generators. These constructive proofs collectively reveal that for these physically important states, the scaling behavior of $C_{\mathrm{opt}}(\rho)$ is indeed closely related to that of $C_{R}(\rho)$, indicating that $C_{R}(\rho)$ is not merely a theoretical baseline, but a physical quantity with strong predictive power for real computational resources.
@@ G.1 Technical Conventions and Parameterization
To ensure the rigor of the arguments, we first unify the technical conventions of this appendix. All logarithms $\log$ are taken to base $2$. The error metric used is the trace distance, $(1/2)\cdot||.||_{1}$. All core physical quantities in this paper, such as Reference-Contingent Complexity $C_{R}(\rho)$, the reference state $\sigma_{R}$, the subspace dimension $d_{R}$, and the information bandwidth parameter $\Gamma_{R}$, follow the definitions in Chapter 2 of the main text. According to the Stinespring dilation theorem [64], any completely positive and trace-preserving (CPTP) map can be realized by a unitary evolution on a larger Hilbert space, followed by tracing out an ancillary system. This fundamental principle ensures that all seemingly non-unitary constructions in this appendix (such as measurements, random mixing) can be compiled into an equivalent, post-selection-free quantum circuit, with the constant factors in their costs being uniformly absorbed into the final upper bound constants. A crucial clarification concerns the parameter dependence of the upper bound constants. To avoid circular logic, we must precisely define the scope of independence for these constants.
@@
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\leq a(\Gamma_{R})\cdot C_{R}(\rho)+b(\Gamma_{R})\cdot(\log d_{R}+\log(1/\varepsilon))$ (G.1)
@@ G.2 Case One: Efficiently Samplable Classical Diagonal States
Assumption G.1 (Efficient Amplitude Preparation). A probability distribution $p$ is considered efficiently preparable in amplitude if there exists a unitary circuit $U_{\mathrm{amp}}$, acting on $n$ ancillary qubits, that can prepare its amplitude-encoded state $|\psi_{p}\rangle=\sum_{x}\sqrt{p_{x}}|x\rangle$ within a precision $\varepsilon^{\prime}$ in a number of atomic instruction steps bounded by $\mathrm{len}(U_{\mathrm{amp}})\leq c_{1}\cdot n+c_{2}\cdot\log(1/\varepsilon^{\prime})$. This assumption covers an important class of probability distributions in physics and algorithms, for which sufficient conditions include, but are not limited to: (i) the distribution can be described by an autoregressive model, and its conditional probabilities $p(x_{i}|x_{<i})$ can be efficiently computed classically; (ii) the distribution comes from a probabilistic graphical model with low tree-width [65, 66, 67, 68].
@@
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\leq b(\Gamma_{R})\cdot(\log d_{R}+\log(1/\varepsilon))$ (G.2)
@@
The entire process is a deterministic CPTP map, and the total upper bound on the number of atomic instruction steps is $O(\log d_{R})+O(\log(1/\varepsilon))$, which proves Eq. (G.2). This result is in high agreement with Conjecture 6.1. $C_{R}(\rho)$ is determined by $(\log d_{R}-H(p))$, while the upper bound on $C_{\mathrm{opt}}(\rho)$ is dominated by $\log d_{R}$. The scaling behavior of the two is consistent when $H(p)$ is much smaller than $\log d_{R}$, and when $H(p)$ approaches $\log d_{R}$, the term $b\cdot\log d_{R}$ provides the correct "fallback" upper bound.
@@ G.3 Case Two: Low-Rank, High-Occupation States and Hierarchical Projection Construction
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\leq a(\Gamma_{R})\cdot C_{R}(\rho)+b(\Gamma_{R})\cdot(\log d_{R}+\log(1/\varepsilon))$ (G.3)
@@
Proof of Lemma. We can construct $T_{\Pi}$ by concatenating $m$ "measure-reprepare-depolarize" quantum instruments $\Phi^{(j)}$. For each $P_{j}$, define $\Phi^{(j)}(X)=P_{j}\cdot X\cdot P_{j}+\mathrm{Tr}[(I-P_{j})\cdot X]\cdot(P_{j}/\mathrm{Tr}(P_{j}))$. This is a post-selection-free CPTP map, which physically corresponds to measuring the eigenvalue of $P_{j}$: if the outcome is +1+1, the state is kept; if the outcome is 1-1, that part of the state is reinjected into the +1+1 subspace of $P_{j}$. Since all $P_{j}$ commute, these $m$ channels can be applied sequentially, with each step compressing the state into a smaller subspace, ultimately mapping $\sigma_{R}$ exactly to $\Pi/r$. The implementation cost of each step $\Phi^{(j)}$ depends on the structure of $P_{j}$. If $P_{j}$ acts on a local subsystem, this operation can be realized via a local $2$-design with $O(1)$ cost [69, 70]. If $P_{j}$ is non-local, its implementation cost is absorbed as a constant factor into the total cost of $O(\log(d_{R}/r))$. Using this lemma, we can prepare $\rho$ by introducing a control qubit: with probability $p$, apply the channel $T_{\Pi}$ to $\sigma_{R}$, and with probability 1p1-p, prepare the background state $\tau$. The total cost is dominated by $O(\log(d_{R}/r))$ and the cost of preparing $\tau$, with an upper bound of $O(\log d_{R})$. On the other hand, from the convexity of relative entropy, the leading term of $C_{R}(\rho)$ is determined by $p\cdot\log(d_{R}/r)$. Therefore, the upper bound on $C_{\mathrm{opt}}(\rho)$ is consistent in scaling with the leading term of $C_{R}(\rho)$, thus verifying Eq. (G.3).
@@ G.4 Case Three: Sparse Graph States and Efficiently Encodable Logical Pure States
Definition G.4 (Family of Efficiently Encodable Logical Pure States). A logical pure state $\rho=|\psi_{L}\rangle\langle\psi_{L}|$ is a pure state located within a $d_{R}=2^{k}$-dimensional code subspace $H_{R}$. It is considered efficiently encodable if there exists an encoding unitary circuit $U_{\mathrm{enc}}$ that can map a simple product state $|s_{0}\rangle$ to a logical ground state $|0\rangle_{L}$ within the code space, in a number of atomic instruction steps bounded by $\mathrm{len}(U_{\mathrm{enc}})\leq c\cdot k$. Logical operations can then be performed on this state. Typical representatives of this class of states are: (i) graph states with bounded vertex degree ($|E|=O(k)$); (ii) logical states of stabilizer codes that are geometrically sparse or satisfy LDPC conditions [15, 71].
@@
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\leq a(\Gamma_{R})\cdot C_{R}(\rho)+b(\Gamma_{R})\cdot\log(1/\varepsilon)$ (G.4)
@@
$C_{R}(\rho)=\frac{\log d_{R}-S(\rho)}{\log\Gamma_{R}}=\frac{k}{\log\Gamma_{R}}$ (G.5)
Comparing the upper bound on $C_{\mathrm{opt}}(\rho)$, which is $O(k)$, with $C_{R}(\rho)=k/\log\Gamma_{R}$, we find that they differ only by a factor of $a(\Gamma_{R})=O(\log\Gamma_{R})$, and their scaling behaviors are identical. This directly proves Eq. (G.4), providing the strongest corroborating evidence for Conjecture 6.1. Conclusion: Through a constructive analysis of three families of states that are physically and computationally distinct but are all "well-structured," we have consistently found that the upper bound on their optimal preparation complexity, $C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$, follows the form $C_{\mathrm{opt}}\leq a(\Gamma_{R})\cdot C_{R}(\rho)+b(\Gamma_{R})\cdot(\log d_{R}+\log(1/\varepsilon))$. This series of non-trivial pieces of evidence strongly supports the validity of Conjecture 6.1 proposed in the main text. This profoundly suggests that Reference-Contingent Complexity $C_{R}(\rho)$ is not just an abstract, possibly loose, theoretical lower bound, but is, for a broad class of physically relevant states, a "near-tight" physical benchmark that can effectively estimate the true computational complexity.
@@ H.4 Model Parameters
Atomic Instruction Alphabet. The set composed of all possible atomic instructions $(u,s)$, whose size is $\Gamma_{R}$.
@@ Appendix I Complete Derivation of the “Work–Complexity” and “Time–Complexity” Relations
This appendix aims to provide a complete and rigorous mathematical derivation and boundary condition analysis for the “Work–Complexity” identity (Proposition 5.3) and the “Process Time Lower Bound” (Theorem 5.2) proposed in Sec. 5.4 of the main text. To ensure the rigor of the discussion and consistency of the framework, we first clarify the following conventions: in this appendix, unless explicitly specified as $\log_{2}$, all instances of $\log$ refer to the natural logarithm $\ln$, and entropy is measured in nats, so as to maintain consistency with the thermodynamic framework. When citing content from the main text, a standard unit conversion between $\mathrm{bits}\leftrightarrow\mathrm{nats}$ will be assumed[35][72].

### v3

@@ Appendix F Generalized Settings and Framework Robustness
This appendix aims to explore the applicability of the Reference-Contingent Complexity (RCC) framework in more generalized settings and to address profound questions regarding the objectivity of its core constructs. Section F.1 generalizes the theory to reference sets defined by non-Abelian symmetries. Section F.2 provides a rigorous scheme for a conservative lower bound for handling systems that require an energy or mode cutoff (such as quantum field theories). Finally, Section F.3 provides physical, objectivization criteria for the selection of the reference set and proves the robustness of the framework with respect to this choice[CasiniHuerta2009, BartlettRudolphSpekkens2007].
@@ F.1 Non-Abelian Reference Sets
$\Pi_{R}=\bigoplus_{\lambda\in\Lambda}(I_{V_{\lambda}}\otimes Q_{\lambda})\\ d_{R}=\mathrm{Tr}(\Pi_{R})=\sum_{\lambda\in\Lambda}d_{\lambda}\cdot r_{\lambda}\\ \sigma_{R}=\Pi_{R}/d_{R}$ (85)
@@ F.2 Conservative Lower Bound under Energy or Mode Cutoff
$D(\rho_{\mathrm{phys}}||\tilde{\sigma}_{R}^{(\Lambda)})\geq D_{\mathrm{Bern}}(q||1-\delta)+q\cdot D(\rho^{(\Lambda)}||\sigma_{R}^{(\Lambda)})$ (88)
@@
$D(\rho_{\mathrm{phys}}||\tilde{\sigma}_{R}^{(\Lambda)})\geq q\cdot D(\rho^{(\Lambda)}||\sigma_{R}^{(\Lambda)})$ (89)
@@ F.3 Acceptability and Objectivity of the Reference Set
$\bigl|C_{R}(\rho)-C_{R^{\prime}}(\rho)\bigr|$ $\leq\biggl|\frac{\log(d_{R})}{\log(\Gamma_{R})}-\frac{\log(d_{R^{\prime}})}{\log(\Gamma_{R^{\prime}})}\biggr|$ (90)
$\quad+\biggl|\frac{1}{\log(\Gamma_{R})}-\frac{1}{\log(\Gamma_{R^{\prime}})}\biggr|\,S(\rho)$
@@
$0\leq C_{R}(\rho)-C_{R^{\prime}}(\rho)=\frac{\log(d_{R})-\log(d_{R^{\prime}})}{\log(\Gamma_{R})}$ (91)
@@ Appendix G Near-Tightness of Reference-Contingent Complexity
This appendix aims to provide constructive, non-trivial, positive evidence for Conjecture 6.1 (the near-tightness of $RCC$), which was proposed in Section VI.2 of the main text. The conjecture posits that Reference-Contingent Complexity $C_{R}(\rho)$ is not only a lower bound on the universal circuit complexity $C_{\mathrm{opt}}(\rho)$, but that for a class of well-structured families of states, it can also be bounded from above by $C_{R}(\rho)$. We will systematically corroborate this conjecture by constructing complexity upper bounds for three core families of states that are widely present in quantum information and physics: efficiently samplable classical diagonal states, low-rank states constructible via hierarchical projections, and efficiently encodable logical pure states with sparse generators. These constructive proofs collectively reveal that for these physically important states, the scaling behavior of $C_{\mathrm{opt}}(\rho)$ is indeed closely related to that of $C_{R}(\rho)$, indicating that $C_{R}(\rho)$ is not merely a theoretical baseline, but a physical quantity with strong predictive power for real computational resources.
@@ G.1 Technical Conventions and Parameterization
To ensure the rigor of the arguments, we first unify the technical conventions of this appendix. All logarithms $\log$ are taken to base $2$. The error metric used is the trace distance, $(1/2)\cdot||.||_{1}$. All core physical quantities in this paper, such as Reference-Contingent Complexity $C_{R}(\rho)$, the reference state $\sigma_{R}$, the subspace dimension $d_{R}$, and the information bandwidth parameter $\Gamma_{R}$, follow the definitions Sec. 2 of the main text. According to the Stinespring dilation theorem [Stinespring1955], any completely positive and trace-preserving (CPTP) map can be realized by a unitary evolution on a larger Hilbert space, followed by tracing out an ancillary system. This fundamental principle ensures that all seemingly non-unitary constructions in this appendix (such as measurements, random mixing) can be compiled into an equivalent, post-selection-free quantum circuit, with the constant factors in their costs being uniformly absorbed into the final upper bound constants. A crucial clarification concerns the parameter dependence of the upper bound constants. To avoid circular logic, we must precisely define the scope of independence for these constants.
@@
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\leq a(\Gamma_{R})\cdot C_{R}(\rho)+b(\Gamma_{R})\cdot(\log d_{R}+\log(1/\varepsilon))$ (92)
@@ G.2 Case One: Efficiently Samplable Classical Diagonal States
Assumption G.1 (Efficient Amplitude Preparation). A probability distribution $p$ is considered efficiently preparable in amplitude if there exists a unitary circuit $U_{\mathrm{amp}}$, acting on $n$ ancillary qubits, that can prepare its amplitude-encoded state $|\psi_{p}\rangle=\sum_{x}\sqrt{p_{x}}|x\rangle$ within a precision $\varepsilon^{\prime}$ in a number of atomic instruction steps bounded by $\mathrm{len}(U_{\mathrm{amp}})\leq c_{1}\cdot n+c_{2}\cdot\log(1/\varepsilon^{\prime})$. This assumption covers an important class of probability distributions in physics and algorithms, for which sufficient conditions include, but are not limited to: (i) the distribution can be described by an autoregressive model, and its conditional probabilities $p(x_{i}|x_{<i})$ can be efficiently computed classically; (ii) the distribution comes from a probabilistic graphical model with low tree-width [LarochelleMurray2011, Germain2015MADE, VanDenOord2016PixelCNN, KollerFriedman2009].
@@
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\leq b(\Gamma_{R})\cdot(\log d_{R}+\log(1/\varepsilon))$ (93)
@@
The entire process is a deterministic CPTP map, and the total upper bound on the number of atomic instruction steps is $O(\log d_{R})+O(\log(1/\varepsilon))$, which proves Eq. (93). This result is in high agreement with Conjecture 6.1. $C_{R}(\rho)$ is determined by $(\log d_{R}-H(p))$, while the upper bound on $C_{\mathrm{opt}}(\rho)$ is dominated by $\log d_{R}$. The scaling behavior of the two is consistent when $H(p)$ is much smaller than $\log d_{R}$, and when $H(p)$ approaches $\log d_{R}$, the term $b\cdot\log d_{R}$ provides the correct "fallback" upper bound.
@@ G.3 Case Two: Low-Rank, High-Occupation States and Hierarchical Projection Construction
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\leq a(\Gamma_{R})\cdot C_{R}(\rho)+b(\Gamma_{R})\cdot(\log d_{R}+\log(1/\varepsilon))$ (94)
@@
Proof of Lemma. We can construct $T_{\Pi}$ by concatenating $m$ "measure-reprepare-depolarize" quantum instruments $\Phi^{(j)}$. For each $P_{j}$, define $\Phi^{(j)}(X)=P_{j}\cdot X\cdot P_{j}+\mathrm{Tr}[(I-P_{j})\cdot X]\cdot(P_{j}/\mathrm{Tr}(P_{j}))$. This is a post-selection-free CPTP map, which physically corresponds to measuring the eigenvalue of $P_{j}$: if the outcome is +1+1, the state is kept; if the outcome is 1-1, that part of the state is reinjected into the +1+1 subspace of $P_{j}$. Since all $P_{j}$ commute, these $m$ channels can be applied sequentially, with each step compressing the state into a smaller subspace, ultimately mapping $\sigma_{R}$ exactly to $\Pi/r$. The implementation cost of each step $\Phi^{(j)}$ depends on the structure of $P_{j}$. If $P_{j}$ acts on a local subsystem, this operation can be realized via a local $2$-design with $O(1)$ cost [BrandaoHarrowHorodecki2016, HarrowLow2009]. If $P_{j}$ is non-local, its implementation cost is absorbed as a constant factor into the total cost of $O(\log(d_{R}/r))$. Using this lemma, we can prepare $\rho$ by introducing a control qubit: with probability $p$, apply the channel $T_{\Pi}$ to $\sigma_{R}$, and with probability 1p1-p, prepare the background state $\tau$. The total cost is dominated by $O(\log(d_{R}/r))$ and the cost of preparing $\tau$, with an upper bound of $O(\log d_{R})$. On the other hand, from the convexity of relative entropy, the leading term of $C_{R}(\rho)$ is determined by $p\cdot\log(d_{R}/r)$. Therefore, the upper bound on $C_{\mathrm{opt}}(\rho)$ is consistent in scaling with the leading term of $C_{R}(\rho)$, thus verifying Eq. (94).
@@ G.4 Case Three: Sparse Graph States and Efficiently Encodable Logical Pure States
Definition G.4 (Family of Efficiently Encodable Logical Pure States). A logical pure state $\rho=|\psi_{L}\rangle\langle\psi_{L}|$ is a pure state located within a $d_{R}=2^{k}$-dimensional code subspace $H_{R}$. It is considered efficiently encodable if there exists an encoding unitary circuit $U_{\mathrm{enc}}$ that can map a simple product state $|s_{0}\rangle$ to a logical ground state $|0\rangle_{L}$ within the code space, in a number of atomic instruction steps bounded by $\mathrm{len}(U_{\mathrm{enc}})\leq c\cdot k$. Logical operations can then be performed on this state. Typical representatives of this class of states are: (i) graph states with bounded vertex degree ($|E|=O(k)$); (ii) logical states of stabilizer codes that are geometrically sparse or satisfy LDPC conditions [Fowler2012, Kitaev2003].
@@
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\leq a(\Gamma_{R})\cdot C_{R}(\rho)+b(\Gamma_{R})\cdot\log(1/\varepsilon)$ (95)
@@
$C_{R}(\rho)=\frac{\log d_{R}-S(\rho)}{\log\Gamma_{R}}=\frac{k}{\log\Gamma_{R}}$ (96)
Comparing the upper bound on $C_{\mathrm{opt}}(\rho)$, which is $O(k)$, with $C_{R}(\rho)=k/\log\Gamma_{R}$, we find that they differ only by a factor of $a(\Gamma_{R})=O(\log\Gamma_{R})$, and their scaling behaviors are identical. This directly proves Eq. (95), providing the strongest corroborating evidence for Conjecture 6.1. Conclusion: Through a constructive analysis of three families of states that are physically and computationally distinct but are all "well-structured," we have consistently found that the upper bound on their optimal preparation complexity, $C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$, follows the form $C_{\mathrm{opt}}\leq a(\Gamma_{R})\cdot C_{R}(\rho)+b(\Gamma_{R})\cdot(\log d_{R}+\log(1/\varepsilon))$. This series of non-trivial pieces of evidence strongly supports the validity of Conjecture 6.1 proposed in the main text. This profoundly suggests that Reference-Contingent Complexity $C_{R}(\rho)$ is not just an abstract, possibly loose, theoretical lower bound, but is, for a broad class of physically relevant states, a "near-tight" physical benchmark that can effectively estimate the true computational complexity.
@@ H.4 Model Parameters
Atomic Instruction Alphabet.
The set composed of all possible atomic instructions $(u,s)$, whose size is $\Gamma_{R}$.
Structon
@@ Appendix I Complete Derivation of the “Work–Complexity” and “Time–Complexity” Relations
This appendix aims to provide a complete and rigorous mathematical derivation and boundary condition analysis for the “Work–Complexity” identity Proposition 5.3 and the “Process Time Lower Bound” Theorem 5.2 proposed in V.4 of the main text. To ensure the rigor of the discussion and consistency of the framework, we first clarify the following conventions: in this appendix, unless explicitly specified as $\log_{2}$, all instances of $\log$ refer to the natural logarithm $\ln$, and entropy is measured in nats, so as to maintain consistency with the thermodynamic framework. When citing content from the main text, a standard unit conversion between $\mathrm{bits}\leftrightarrow\mathrm{nats}$ will be assumed[liu2025heuristic][zou2023information].